OpenStudy (anonymous):

1 to answer get a medal 5n-16-8n=-10

OpenStudy (anonymous):

-2

jimthompson5910 (jim_thompson5910):

5n-16-8n=-10 -3n-16=-10 -3n=-10+16 -3n=6 n=6/(-3) n=-2